In which case (in case the air flow flows to the right or left with the constant velocity of U), the drag force acting on the object becomes high? Please explain.

U? U?
Answer #1

As the Drag Force,Fd is given by :

FD=(Cd\times A\times \rho \times V^{2})\div 2

where :

Cd = Drag Coefficient

A = Reference Area

Rho = Fluid Density

V = Velocity of fluid

Now here Velocity U, is same from both sides,Density is also same for fluid, Cd is also constant

So only Area is left which is proportional to the Drag Force. So from the figure Right side has less frontal area than the left one so Drag force on the left side is more.